The environmental protection agency publishes figures on solid waste generation in a city. one of the year, the average number of waste generated per person per day was 3.59 pounds. Suppose the daily amount of waste generated per person is normally distributed with a standard deviation of 1.04 pounds. What is the amount of waste generated per person by 67.72% of the people in the city?


1
Expert's answer
2021-02-24T12:48:05-0500

We have that

"\mu=3.59"

"\sigma=1.04"

"P(X<x) = 0.6772 \implies Z=0.46"


"Z=\frac{x-\mu}{\sigma}=\frac{x-3.59}{1.04}=0.46\ \implies x=1.04\cdot0.46+3.59=4.0684"

Answer: 4.07